OpenStudy (anonymous):

Solve for x: 4x + 2 = x – 7

OpenStudy (anonymous):

These are really quite simple. Do you have absolutely no clue where to start? Start by shifting all the x to one side.

OpenStudy (anonymous):

basically your just moving this over 4x+2=x-7 -x -x _______________ like such and just keep moving things over to get x on one side and the number on the other

OpenStudy (anonymous):

do you understand

OpenStudy (anonymous):

No , i succck at mathh c;

OpenStudy (anonymous):

You want to get x on only one side of the equation. You want to remove it from the right side, so you subtract it from there. However, to keep both sides equal, you must subtract x on the left side too. So: 3x + 2 = -7 How would you proceed from there, following the same logic?

OpenStudy (anonymous):

@aussy123 , missed out a -ve sign for the 7 there. :)

OpenStudy (anonymous):

@msohcw thx I really didnt notice :)

OpenStudy (anonymous):

soo illl repeat ok ill show you the beginning you simplify 4x+2=x-7 -x -x ____________ 3x +2=- 7 -2 -2 ________________ 3x=-9 now simplify

OpenStudy (anonymous):

if its - 2 isnt it supposed to be - 5 ?

OpenStudy (anonymous):

when you move the 2, you have to use its opposite which is -2

OpenStudy (anonymous):

so it would be , like minus 2 on both sides ?

OpenStudy (anonymous):

You're effectively subtracting 2 from both sides.

OpenStudy (anonymous):

Yep. So that you can keep the two sides equal.

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now at the bottom its minus 3 as well ?

OpenStudy (anonymous):

s

OpenStudy (anonymous):

at the bottom you do the opposite of multiplication which is division and divide 3 over to the -9...and -9 divided by 3 is

OpenStudy (anonymous):

-3 right because ; a - divided by a postive is a negative o;?

OpenStudy (anonymous):

CORRECT!!!

OpenStudy (anonymous):

DO you understand now?

OpenStudy (anonymous):

yeah , i just need you too explain . How too pellet all the x's too one side . Like where do the other numbers go ?

OpenStudy (anonymous):

*put >.<

OpenStudy (anonymous):

ok above when you moved the x over by subtracting it, you ended up with 3x+2=-7 all you had to do is do the same thing to the 2 that you did with the x..and make sure you when you do problems like this to take the easiest road or it will become tricky...but I believe you can do it!!

OpenStudy (anonymous):

Help me start off this one , and ill do the rest c: ? Solve for x: 2x – 1 = x – 6

OpenStudy (anonymous):

Start by shifting the x s again all to one side. How do you think you should do that?

OpenStudy (anonymous):

i think i have the answer xD

OpenStudy (anonymous):

Good job. What grade are you in? Do your teachers not teach you anything >.< ?

OpenStudy (anonymous):

isss it -5 o; ? & well im in summer school thats why ,

OpenStudy (anonymous):

were just on the computers

OpenStudy (anonymous):

Yep. -5 indeed. To check, you can put it back into the equation to see if it works. 2x – 1 = x – 6 2(-5) -1 = (-5) -6 -10-1 = -11 -11 = -11 -11 does equal -11 (of course), so it's right. What's summer school? An advanced program? (I'm not from the US.)

OpenStudy (anonymous):

itss to make up high school credits , so you can go to the next grade c:

OpenStudy (anonymous):

I see So you're 15/16? Hope you're understanding algebra better now. If you've got an iPad/iPhone, go buy DragonBox, it's a great (and ACTUALLY fun) app to learn algebra with.

OpenStudy (anonymous):

yeaah im 15